A college awarded 26 varsity letters in crew, 15 in swimming, and 16 in soccer. If awards went to 46 students and only 2 lettered in all sports, how many students lettered in two of the three sports
Answer by richwmiller(17219)

This is illustrated with a venn diagram.
26+15+16=57
but 2 are in all three.
Remove 4
57-4=53
53-46=7
There are 7 students who lettered in two of three sports.